# RobustBlack: Challenging Black-Box Adversarial Attacks on State-of-the-Art Defenses

### v1

@@ 3 Success Rate of Black-Box Attacks Against Adversarial Training
As preliminaries, we investigate the effectiveness of 12 black-box adversarial attacks against Madry adversarial training (Madry et al., 2017) on a single, relatively simple ResNet50 model (He et al., 2016). The purpose of these experiments is to check to what extent black-box attacks, typically successful against vanilla models, remain as successful when confronted to adversarially trained models. This preliminary experiment provides us with the first insight into the impact of defenses studied in white-box settings on the effectiveness of black-box attacks.
3.1 Experimental Setup
@@
We evaluated the attack success rate of 11 black-box adversarial attacks against a standardly trained model (vanilla) and an adversarially trained model. We selected a ResNet50 vanilla target and a ResNet50 robust target that underwent adversarial training. We opted for the same model architecture to eliminate the variations that could arise from differing model architectures. We followed the Robustbench evaluation protocol (Croce et al., 2020) and reused the original implementation of each attack. All attacks were untargeted and bound to an $L_{\infty}$ maximum perturbation with a distance of $4/255$. We evaluated the ASR on 5000 examples from the ImageNet validation set, we used the same image identifiers (IDs) as in RobustBench. All the experiments are repeated with three randoms seeds.
@@
As outlined in 2.1, we selected nine transfer-based and two query-based adversarial attacks to provide a comprehensive evaluation of adversarial techniques, balancing foundational, and SOTA approaches. For transfer attacks, our selection emphasizes the diversity of techniques by including methods targeting input transformations (TI, DI), gradient optimization (MI, VNI, VMI, ADMIX), universality (UAP), architectural nuances (GHOST), and geometry (LGV). For query-based attacks, we focused on methods that improve query efficiency and performance by leveraging surrogate models, and that rely on different optimization techniques. BASES uses a small dimensional search space by modifying the weights given to each surrogate based on the feedback received from querying the target, while TREMBA learns a low-dimensional embedding then performs an efficient search within this space. As discussed in 2.1, all query attacks in this study require surrogate models. The hyperparameters for all transfer and query attacks are detailed in the Appendix A and B, respectively.
@@
We employed surrogate models that are either individually larger than the target model or collectively larger as an ensemble, allowing us to leverage the full capacity of the attack. For all single-surrogate attacks, including the five FGSM-based baseline attacks, ADMIX, UAP, GHOST, and LGV, the surrogate is a WideResNet50-2 (Zagoruyko, 2016) with standard training. For TREMBA, the surrogates are an ensemble of four models with standard training: VGG16 (Simonyan & Zisserman, 2014), ResNet18, SqueezeNet (Iandola et al., 2016), and GoogLeNet (Szegedy et al., 2015) as provided in their repository. For BASES, the surrogates are an ensemble of 10 models with standard training: VGG16BN, ResNet18, SqueezeNet11, GoogLeNet, MnasNet10 (Tan et al., 2019), DenseNet161 (Huang et al., 2017), EfficientNetB0 (Tan & Le, 2019), RegNetY400MF (Radosavovic et al., 2020), ResNeXt10132x8d (Xie et al., 2017), ConvNeXtSmall (Liu et al., 2022) from . The models used in this study were obtained from the torchvision library (maintainers & contributors, 2016) and the RobustBench benchmark Croce et al. (2020). We excluded the SGM attack from this section because, at the time of this study, its implementation supports only ResNet and DenseNet surrogates, while our setup uses WideResNet50-2.
3.2 Results
@@
We evaluate the ASR of the black-box attacks against standard and adversarially trained Resnet-50 models. We illustrate the mean ASR using bars, and the standard deviation (STD) as error bars in Figure 1. Against the undefended model, all attacks except UAP and LGV achieve more than 40% ASR, with the most recent attacks BASES and TREMBA reaching 81.08% ± 0.49 and 89.56% ± 0.09, respectively. However, all black-box attacks fail against the robustified model with an ASR of less than 4%. The most recent attacks only bring about marginal improvement over the simplest FGSM attacks. The most effective attack in our evaluation, TREMBA achieves 3.56% ASR while TI-FGSM reaches 2.81% ASR. To ensure that the reduced effectiveness of black-box attacks against the adversarially trained ResNet50 model is not due to budget limitation, we quadruple the budget of the best black-box attacks (BASES and TREMBA). Even with this increased budget, both attacks still significantly underperform against the adversarially trained ResNet50 compared to their performance on the vanilla ResNet50 model, as detailed in Appendix C.
@@
Figure 1: The blue bars (with error bars) show the success rates and standard deviations for the vanilla ResNet50 model, while the orange bars (with error bars) show the results for the robust ResNet50 model.
@@
Insight 1 The effectiveness of black-box attacks drastically decreases against a simple adversarially trained model, with the strongest attack seeing its ASR drops from 89.56% to 3.56%.
4 Effectiveness of White-Box Defenses Against Black-Box Attacks
Given that even the most recent advances in black-box attacks remain ineffective against Madry adversarial training, we investigate whether recent innovations in defense mechanisms (studied in white-box settings) further reduce the effectiveness of black-box attacks. In particular, we want to observe whether these defenses designed against white-box attacks (viz. AutoAttack) and established benchmarks (viz. Robustbench), generalize to black-box attacks. A positive answer to these questions would maintain (and even raise) the motivation to develop defenses in white-box settings, whereas a negative answer would raise an urgent necessity of investigating specific defenses against black-box attacks.
4.1 Experimental Setup
We followed the same protocol as in the previous section, in terms of datasets, and adversarial black-box attacks. We selected nine robust target models from the Robustbench leaderboard. We ensured to have an even distribution across the robustness spectrum for the nine targets. The robust accuracy of the most robust model, a Swin-L (Liu et al., 2024) is 59.64% (that is, a success rate of AutoAttack of 25.97%), and the robust accuracy of the least robust model, a ResNet18 (Salman et al., 2020) is 25.32% (for a success rate of AutoAttack of 50.53%). Table 1 presents the architecture and the number of parameters of the models. These nine models have been defended using different categories of mechanisms, listed in Table 2.
We kept the same surrogate ensemble for BASES and TREMBA, and we used the surrogate model ResNet50 with standard training for single-surrogate attacks. In addition to the selected black-box attacks, we incorporated the SGM attack due to its compatibility with the ResNet50 surrogate model.
Table 1: Summary of different models from the RobustBench leaderboard (Croce et al., 2020) and their configurations. Legend of the defenses in Table 2.
@@ 4.1 Experimental Setup
1 Liu et al. (2024) Swin-L (Liu et al., 2021) 59.56 % A Transformer 196.53M
2 Bai et al. (2024) ConvNeXtV2+Swin-L (Liu et al., 2022; 2021) 58.50 % A, B Transformer & Convolution 394.49M
3 Liu et al. (2024) ConvNeXt-L (Liu et al., 2022) 58.48 % A Convolution 197.77M
5 Liu et al. (2024) Swin-B (Liu et al., 2021) 56.16 % A Transformer 87.77M
7 Liu et al. (2024) ConvNeXt-B (Liu et al., 2022) 55.82 % A Convolution 88.59M
12 Singh et al. (2024) ViT-S+ConvStem (Dosovitskiy et al., 2020; Xiao et al., 2021) 48.08 % C Transformer 22.78M
17 Salman et al. (2020) WideResNet50.2 (Zagoruyko, 2016) 38.14 % D Convolution 68.88M
18 Salman et al. (2020) Resnet50 (He et al., 2016) 34.96 % D Convolution 25.56M
21 Salman et al. (2020) Resnet18 (He et al., 2016) 25.32 % D Convolution 11.69M
Table 2: List of the defense mechanisms.
Label Description
A Adv training with large data augmentation, regularisation, weights averaging, pretraining
B Non linear ensemble of two base models robust and vanilla
C Downsample convolutional layers before subsequent network layers
D Standard adversarial training
4.2 Results
@@
We analyze the performance of defenses against black-box attacks by examining the impact of different defense strategies, size of models, and architecture of models, as well as the relationship between AutoAttack success rates and black-box attacks success rates.
Impact of defense mechanism.
We categorize the defended models from Table 1 into four distinct families (A, B, C, and D), reflecting the various training strategies employed to enhance model robustness. In Figure 2 we show the ASR of the 12 black-box attacks against the four defenses (robust models). The models are sorted according to the Robustbench leaderboard, with stronger defenses against AutoAttack positioned to the right. Models using the same category of defense mechanisms but differing architectures are grouped in the same color. We observe that stronger defenses against AutoAttack lead to stronger defenses against black-box attacks. For instance, Defense A (red) outperforms Defense C (yellow), which in turn surpasses Defense D (blue) across all black-box attacks. An exception to this trend is observed with the ensemble Conv+Swin Defense, ranked #2 on Robustbench and mixing mechanisms of types A and B. Although ranked second against AutoAttack, it performs similarly to Defense D when subjected to black-box attacks. This is explained by the composition of defense mechanisms A and B, which combines the adversarially trained model from Defense A with a standard (vanilla) model, as explained in Section 2.2. This hybrid nature introduces a vulnerability: while this defense ranks among the best defenses against AutoAttack, the inclusion of a standard model component causes its performance against black-box attacks to align with that of weaker AutoAttack defenses. This vulnerability arises because black-box attacks often exploit vanilla surrogates, making the defense susceptible due to the standard model’s presence.
Table 3: Success rate on Resnet models trained by Salman et al. (2020).
Attacks Salman ResNet18 Salman ResNet50 Salman WideResNet50.2
AUTOATTACK 50.53 43.84 42.87
@@
Impact of model size.
We compare in Table 3 the success rates of the black-box attacks against similar adversarially trained models. These models do not benefit from advanced defense mechanisms (e.g., synthetic data augmentation, …) and only differ in size. While larger models with adversarial training tend to achieve higher robustness to AutoAttack, the increase in size does not lead to lower success rates. Similar insights are revealed in Table 4, where we compare the robustness of two sizes of ConvNext models and SWIN Transformers. Larger models do not consistently demonstrate stronger robustness to adversarial perturbations in black-box settings, although increased size improves robustness against AutoAttack.
Impact of model architecture.
Liu et al. (2024) demonstrate that the best robustness against AutoAttack can be achieved using indiscriminately Convnets or Transformers models. Our results in Table 4 confirm that for a fixed size, both architectures demonstrated similar robustness against black-box attacks.
Table 4: Success rate on very large models trained by Liu et al. (2024).
Attacks Liu ConvNeXt-B Liu Swin-B Liu ConvNeXt-L Liu Swin-L
AUTOATTACK 28.01 27.14 25.95 25.98
@@
Relationship between AutoAttack success rate and black-box success rate. As shown in Figure 3, the scatter plot illustrates the relationship between the ASR of AutoAttack and that of 12 black-box attacks. Each data point represents the ASR on the y-axis of a given black-box attack (indicated by a color) against a specific robust model (determined by its AutoAttack success rate on the x-axis).
Although the initial ASR of black-box attacks is low (less than 5%), we observe a clear downward trend as robust models improve their effectiveness against AutoAttack, until the best four robust models (with AutoAttack success rate ranging from 25.48% to 27.14%). This vulnerability stems from the lack of correlation between model size and robustness, as these models all employ defense (A) with diverse sizes and architectures, as shown in Table 1.
@@
Insight 2 SoTA defenses against AutoAttack generally correlate with improved black-box robustness. However, having a large model size does not always guarantee better robustness. Additionally, an ensemble defense that achieves similar robustness to a single defense against AutoAttack may be less effective against black-box attacks, especially those using surrogate models similar to the ensemble’s components.
5 Robust surrogates to Improve Black-Box Attacks
The results from the previous section demonstrated that models robust against AutoAttack remain relevant (to some extent) against black-box attacks. A critical concern then is whether these robust "defenses" could not be used by the attacker to improve its black-box attack. Indeed, most of the black-box require designing a surrogate and can therefore leverage the recent advances in model defense to strengthen their attacks. We hypothesize that these new defenses raise security risks by providing stronger surrogates to attackers.
5.1 Experimental Setup
We adhered to the same protocol as in previous sections, maintaining consistency in datasets and adversarial black-box attacks. We considered various configurations regarding the robustness of both the surrogate and target models. For the target models, we selected highly robust models (Swin-L, ConvNeXtV2+Swin-L, ConvNext-L from (Liu et al., 2024) and (Bai et al., 2024)), a moderately robust model (ResNet18 from (Salman et al., 2020)), and a non-robust model (ResNet50). This way, we can measure the effects of using robust surrogates on targets with top defense (A), a simple defense (D) and no defense.
As for surrogate, we used robust and non-robust models. The selection of robust surrogates was based on choosing the highly ranked robust models that are not used as target and are compatible with the attack methods. Hence, we used the robust surrogate RaWideResNet-101-2 from (Peng et al., 2023) for single surrogate attacks, because it is the best defense in Robustbench that supports skip connections (which are essential for the GHOST attack). For ensemble surrogate attacks, we also added the robust surrogate Swin-B from (Liu et al., 2024), as it is the best robust model available, excluding the target models. For single, non-robust surrogate attacks, we used a vanilla WideResNet-101-2 in order to have the same architecture as in the robust surrogate, ensuring fair comparison. For BASES and TREMBA (who use an ensemble of surrogates), we kept the same ensemble of vanilla models as in our previous experiments. We excluded the SGM attack that does not support the WideResNet-101-2 surrogates used in our setup.
5.2 Results
We present in Figure 4 the success rate of black-box attacks using non-robust surrogates and robust surrogates against a non robust target and robust targets.

### v2

@@ III Success Rate of Black-Box Attacks Against Adversarial Training
As preliminaries, we investigate the effectiveness of 13 black-box adversarial attacks against Madry adversarial training [47] on a single, relatively simple ResNet50 model [31]. The purpose of these experiments is to check to what extent black-box attacks, typically successful against vanilla models, remain as successful when confronted to adversarially trained models. This preliminary experiment provides us with the first insight into the impact of defenses studied in white-box settings on the effectiveness of black-box attacks.
III-A Experimental Setup
@@
We selected a ResNet50 vanilla target and a ResNet50 robust target that underwent adversarial training. We use the same model architecture to eliminate the variations that could arise from differing model architectures. We followed the Robustbench evaluation protocol [15] and reused the original implementation of each attack: All attacks are untargeted and bound to a $L_{\infty}$ maximum perturbation with a distance of $4/255$, unless mentioned otherwise. We evaluated the Attack Success Rate (ASR) on 5000 examples from the ImageNet validation set, and we used the same image identifiers (IDs) as in RobustBench.
@@
We selected nine transfer-based adversarial attacks and four query-based adversarial attacks to provide a comprehensive evaluation of adversarial techniques, balancing foundational approaches and SoTA. For transfer attacks, our selection emphasizes the diversity of techniques by including methods targeting input transformations (TI, DI), gradient optimization (MI, VNI, VMI, ADMIX), universality (UAP), architectural nuances (GHOST), and geometry (LGV). For query-based attacks, we included two families. First, pure query attacks, which rely solely on feedback from the target model such that their performance is not influence by the choice of the surrogate model , such as (SQUARE) and (SIGNFLIP). Second, hybrid query attacks, which improve query efficiency and performance by leveraging surrogate models with optimization strategies to reduce the number of queries. Specifically, (BASES) modifies the weights assigned to each surrogate based on feedback from the target, effectively searching a small-dimensional space, while (TREMBA) learns a low-dimensional embedding and conducts an efficient search within this space. The details and hyperparameters for all attacks are provided in the Appendix -C.
@@
For all single-surrogate attacks, including the five FGSM-based baseline attacks, ADMIX, UAP, GHOST, and LGV, the surrogate is a WideResNet50-2 [81] with standard training. For TREMBA, the surrogates are an ensemble of four models with standard training: VGG16 [60], ResNet18, SqueezeNet [35], and GoogLeNet [64] as provided in their repository. For BASES, the surrogates are an ensemble of 10 models with standard training: VGG16BN, ResNet18, SqueezeNet11, GoogLeNet, MnasNet10 [65], DenseNet161 [32], EfficientNetB0 [66], RegNetY400MF [58], ResNeXt10132x8d [80], ConvNeXtSmall [46] . The models used in this study were obtained from the torchvision library [50] and the RobustBench benchmark [15]. We excluded the SGM attack from this section because, at the time of this study, its implementation supports only ResNet and DenseNet surrogates, while our setup uses WideResNet50-2.
III-B Results
@@
We evaluated the ASR of the black-box attacks against standard and adversarially trained Resnet-50 models. We illustrate the ASR using bars in Figure 1. Under a restrictive perturbation budget of 4/255, all attacks except UAP and SIGNFLIP achieve more than 40% ASR against the undefended model, with the most recent attacks LGV, BASES and TREMBA reaching 90.11% ± 0.12, 81.08% ± 0.49, and 89.56% ± 0.09, respectively. However, all black-box attacks fail against the robustified model with an ASR of less than 17%. To rule out query-budget limitations, we quadruple the budgets of the BASES and TREMBA hybrid attacks. Even with this increase, both attacks still perform far worse against the adversarially trained ResNet50 than against the vanilla model (see Appendix A-A).
When the perturbation budget is increased to 16/255 and the query budget for query-only attacks is doubled, black-box attacks can achieve higher ASR. Under these settings, the performance gap between the undefended and defended models narrows, with the Square Attack reaching an ASR of 79%.
@@
Figure 1: The blue bars show the success rates and standard deviations for the vanilla ResNet50 model, while the orange bars show the results for the robust ResNet50 model. Bars in solid color are with attack distance $\epsilon=4/255$, while dashed bars are with attack distance $\epsilon=16/255$
III-C Parameter Sensitivity Analysis
We study in the following the impact of increasing perturbation and computation budgets. In Fig.2 we present the results for the increasing $\epsilon$ budgets, and report the exact performance with increased number of iterations and queries in Appendix B. Our results confirm that significantly increasing the perturbation budget does not lead to a collapse of robustness of adversarially trained models and that TREMBA is the only attack with a success rate increasing over 30% for extremely large budgets. The results in Table II show that increasing iterations has negligible impact, and Table III that increasing the number of queries has also negligible impact.
Figure 2: Sensitivity study of the impact of the perturbation budget $\epsilon$.
TABLE II: Parameter study of the impact of the iteration budget.
Attack Iteration=N Iteration=2N Iteration=5N
MI-FGSM 88.79 $\rightarrow$ 10.52 91.63 $\rightarrow$ 11.61 93.79 $\rightarrow$ 12.67
DI-FGSM 94.55 $\rightarrow$ 12.11 97.32 $\rightarrow$ 13.30 99.17 $\rightarrow$ 14.38
TI-FGSM 83.53 $\rightarrow$ 15.61 89.31 $\rightarrow$ 17.61 93.45 $\rightarrow$ 18.69
VMI-FGSM 94.79 $\rightarrow$ 12.74 97.73 $\rightarrow$ 14.30 99.06 $\rightarrow$ 14.82
VNI-FGSM 96.94 $\rightarrow$ 13.21 98.62 $\rightarrow$ 14.67 99.45 $\rightarrow$ 14.98
ADMIX 98.28 $\rightarrow$ 14.24 98.90 $\rightarrow$ 14.58 99.43 $\rightarrow$ 15.02
UAP 70.87 $\rightarrow$ 9.93 70.78 $\rightarrow$ 9.93 72.38 $\rightarrow$ 10.02
GHOST 94.40 $\rightarrow$ 11.27 96.48 $\rightarrow$ 12.21 98.41 $\rightarrow$ 13.39
LGV 99.66 $\rightarrow$ 15.16 99.92 $\rightarrow$ 16.61 99.97 $\rightarrow$ 16.39
TABLE III: Parameter study of the impact of number of query budgets.
Model Attack Query=N Query=2N
Vanilla BASES 99.56 99.66
Vanilla TREMBA 99.84 99.90
Robust BASES 13.76 14.05
Robust TREMBA 32.67 37.79
@@ III-C Parameter Sensitivity Analysis
Insight 1 The effectiveness of black-box attacks drastically decreases against a simple adversarially trained model, with the strongest attack (Square) seeing its ASR drop from 79.04% to 16.06%.
IV Effectiveness of White-Box Defenses Against Black-Box Attacks
Our first insight demonstrates that simple Madry adversarial training is an effective defense against most SoTA black-box attacks under a restrictive attack budget. We explore in this section whether improved defenses manage to also improve against SoTA black-box attacks in all the settings. These new defenses have been designed to overcome AutoAttack, the SoTA white-box attack, and we are the first to confront them to SoTA black-box attacks.
IV-A Experimental Setup
We followed the evaluation protocol of the first study on nine robust models from Robustbench. We selected models that were (1) peer-reviewed, (2) compatible with all the attacks (for instance Skip connections to run GHOST attacks), and (3) available in a diverse set of architectures (transformers , convolutions), size (Base and Large variants), and defense mechanisms (ensembling, synthetic data augmentation, downsampling layers). Our selection strategy allows to study the impact of defense mechanisms as a whole, and the impact of size and architecture within the same defense mechanism. We provide more details in Appendix -B.
The most robust model, a Swin-L [43] achieves 59.64% robust accuracy, and the least robust model, a ResNet18 [59] achieves 25.32% robust accuracy. Table IV presents the architecture and the number of parameters of the models.
These nine models have been defended using different mechanisms. Mechanism A involves adversarial training with extensive data augmentation, regularization, weight averaging, and extensive pretraining. Mechanism B involves non linear ensemble of two models, one defended with adversarial training and one vanilla model with standard training. Mechanism C mitigates adversarial examples using downsample convolutional layers before subsequent network layers. All models with standard adversarial training are denoted with defense D.
TABLE IV: Robust models included in our evaluation and their configurations. Defense A: Adv training with large data augmentation, regularization, weights averaging, pretraining; Defense B: Non linear ensemble of two base models robust and vanilla; Defense C: Downsample convolutional layers before subsequent network layers; Defense D: Standard adversarial training
@@ IV-A Experimental Setup
1 [43] Swin-L [45] 59.56 % A Transformer 196.53M
2 [5] ConvNeXtV2+Swin-L [46, 45] 58.50 % B Transformer & Convolution 394.49M
3 [43] ConvNeXt-L [46] 58.48 % A Convolution 197.77M
5 [43] Swin-B [45] 56.16 % A Transformer 87.77M
7 [43] ConvNeXt-B [46] 55.82 % A Convolution 88.59M
12 [61] ViT-S+ConvStem [24, 78] 48.08 % C Transformer 22.78M
17 [59] WideResNet50.2 [81] 38.14 % D Convolution 68.88M
18 [59] Resnet50 [31] 34.96 % D Convolution 25.56M
21 [59] Resnet18 [31] 25.32 % D Convolution 11.69M
IV-B Results
@@
We study the performance of defenses against black-box by considering the impact of different defense strategies, the size of the models, and the architecture of the models. We also analyze these results by discussing the relationship between AutoAttack success rates and black-box attack success rates.
Impact of defense mechanism
We report in Figure 3 the ASR of the 14 black-box attacks against each of the defended models. The models are sorted from left to right according to the Robustbench leaderboard, with stronger defenses against AutoAttack on the right. Models using the same category of defense mechanisms are grouped with the same color. Defenses from the family A (red) outperform all other defenses. Defense C (orange) is better than ensembling (green), and vanilla adversarial training (blue). Surprisingly, the Conv + Swin ensemble model (orange, ranked 2 against AutoAttack) significanly underperforms models supposedly ”weaker” according to Robustbench, and is often weaker than old defenses using only standard training. Conv + Swin is weaker than the vanilla adversarial models against MI-FGSM, DI-FGSM, VM-FGSM, VNI-FGSM, SGM, GHOST, and BASES. We hypothesize that this vulnerability arises because black-box attacks often exploit vanilla surrogates, making the ensemble defense susceptible due to the presence of the standard model.
These trends remain consistent when increasing the epsilon values from $4/255$ to $8/255$ and $16/255$, as well as when extending the iteration budget from N to 2N and 5N. Similarly, increasing the query budget from $N$ to $2N$ does not alter the observed pattern. The detailed results are presented in Appendix B.
TABLE V: Success rate on Resnet models trained by Salman et al.[59].
Attacks ResNet18 ResNet50 WideResNet50.2
AA 50.53 43.84 42.87
@@
SQUARE 20.58±0.21 16.05±0.21 15.09±0.1
SIGNFLIP 4.6±0.06 4.29±0.1 3.79±0.06
Impact of model size and architecture
We focus now on models using the same defense mechanism, and compare the impact of the size of model and archtitectures.
We compare in Table V the success rate of the black-box attacks against models defended with standard adversarial training (family D). These models do not benefit from advanced defense mechanisms (e.g., synthetic data augmentation, …) and only differ in size. While scaling up the models from ResNet18 to WideResNet50 lowers AutoAttack (AA) success from $50.53\%$ to $42.87\%$, this pattern does not hold always for black-box attacks. The black-box attacks that suffer from the increase in target size are TREMBA, SQUARE, and SIGNFLIP with the SQUARE attack dropping from $20.58\%$ to $15.09\%$.
Moving to Defense A trained models (extensive data enhancement, much larger models, and longer trainings), we report the impact of the increase in model size in Table VI. While increasing the size of the model (ConvNeXt-B $\rightarrow$ ConvNeXt-L, Swin-B $\rightarrow$ Swin-L) decreases the ASR of AutoAttack, it has no effect on black-box attacks. SIGNFLIP ASR marginally decreases from $1.63\%$ to $1.42\%$ on convolution models with increased size, but marginally improves from $1.28\%$ to $1.65\%$ on Swin transformers with increased size.
TABLE VI: Success rate on very large models trained by Liu et al.[43].
Attacks ConvNeXt-B Swin-B ConvNeXt-L Swin-L
AA 28.01 27.14 25.95 25.98
@@
SQUARE 8.9±0.1 8.71±0.12 8.56±0.04 8.65±0.11
SIGNFLIP 1.63±0.06 1.28±0.03 1.42±0.03 1.65±0.12
Relationship between AutoAttack success rate and black-box success rate
As shown in Figure 4, the scatter plot illustrates the relationship between the ASR of AutoAttack and that of 14 black-box attacks. Each data point represents the ASR on the y-axis of a given black-box attack (indicated by a color) against a specific robust model (determined by its AutoAttack success rate on the x-axis). We observe a clear downward trend as robust models improve their effectiveness against AutoAttack, until the best four robust models (with AutoAttack success rate ranging from 25.48% to 27.14%). This vulnerability stems from the lack of correlation between model architecture and robustness, as these models all employ defense (A) with diverse sizes and architectures, as shown in Table IV.
@@
Insight 2 SoTA defenses improve robustness to black-box attacks. However, larger models do not guarantee higher robustness and effective ensemble defense against AutoAttack may be less effective against black-box attacks, especially those using surrogate models similar to the ensemble’s components. Large data augmentation is the best defense against black-box attacks.
V Robust surrogates to Improve Black-Box Attacks
The results from the previous section demonstrated that models robust against AutoAttack remain relevant (to some extent) against black-box attacks. A critical concern then is whether these robust ”defenses” could not be used by the attacker to improve its black-box attack. Indeed, most of the black-box require designing a surrogate and can therefore leverage the recent advances in model defense to strengthen their attacks. We hypothesize that these new defenses raise security risks by providing stronger surrogates to attackers.
V-A Experimental Setup
We adhered to the same protocol as in previous sections. We considered various configurations of surrogate and target models. For the target models, we selected highly robust models (Swin-L, ConvNeXtV2+Swin-L, ConvNext-L from [43] and [5]), a moderately robust model (ResNet18 from [59]), and a non-robust model (ResNet50). This way, we can measure the effects of using robust surrogates on targets with top defense (A), a simple defense (D) and no defense.
We used the robust surrogate WideResNet-101-2 from [54] for single surrogate attacks, because it is the best defense in Robustbench that supports skip connections (which are essential for the GHOST attack). For ensemble surrogate attacks, we also added the robust surrogate Swin-B from [43], as it is the best robust model available, excluding target models. For single, non-robust surrogate attacks, we used a vanilla WideResNet-101-2 in order to have the same architecture as in the robust surrogate, ensuring fair comparison. For BASES and TREMBA (which use an ensemble of surrogates), we kept the same ensemble of vanilla models as in our previous experiments. We excluded the SGM attack that does not support the WideResNet-101-2 surrogates used in our setup.
V-B Results
We present in Figure 5 the success rate of black-box attacks using non-robust surrogates and robust surrogates against a non-robust target and robust targets.